OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (0, -5)?

OpenStudy (tkhunny):

The one that looks like this: (y+5)^2 = (x) or (y+5) = (x)^2

OpenStudy (anonymous):

how did you get that?

OpenStudy (xapproachesinfinity):

Where is the equations, he answered based on the general form of a parabola

OpenStudy (tkhunny):

With only one piece of information, the vertex, there is not a unique solution. The Vertex Form get's us that close. Still a little information missing.

OpenStudy (xapproachesinfinity):

i think the question have multiple choice answers